2(-1)+1-(-1-2)=2-[-2(-1)]
1
Expert's answer
2012-09-13T07:18:18-0400
2(-1)+1-(-1-2)=-2+1+3=2

2-[-2(-1)]=2-2=0
2<>0 so that's incorrect